OpenStudy (anonymous):

Solve the system of equations. 3x + 3y = –6 –x + y = 0 A. x = 10; y = –3 B. x = 22; y = 7 C. x = –1; y = –1 D. x = 14; y = 5

OpenStudy (anonymous):

Simplifying 3x + -3y = -6 Solving 3x + -3y = -6 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3y' to each side of the equation. 3x + -3y + 3y = -6 + 3y Combine like terms: -3y + 3y = 0 3x + 0 = -6 + 3y 3x = -6 + 3y Divide each side by '3'. x = -2 + y Simplifying x = -2 + y

OpenStudy (anonymous):

@chiefarnav i don't understand....how did you get that? thats not an option..

OpenStudy (alekos):

No its not

OpenStudy (alekos):

I think he's led you up the garden path

OpenStudy (alekos):

Just multiply the 2nd equation by 3 To get -3x+3y=0

OpenStudy (alekos):

If you then add the two equations you get 6y=-6 or y=-1 If you subtract the equations you get x=-1

OpenStudy (radar):

From the 2nd equation, you know that x = y! Now it becomes pretty simple to solve by substitution.